OpenStudy (kimberly_pr):

Help please??!! Will medal and fan Simplify parenthesis 1 minus cosine theta parenthesis times parenthesis 1 plus cosine theta parenthesis divided by parenthesis 1 minus sine theta parenthesis times parenthesis 1 plus sine theta parenthesis

OpenStudy (kimberly_pr):

Hey @agent0smith idk if you left yet, but if you did have fun!! And if you haven't help please? :/ lol

OpenStudy (kimberly_pr):

Oh and @welshfella can you help please?

OpenStudy (agent0smith):

Sorry, leaving in minutes. Post a screenshot cos no one ever is deciphering "parenthesis 1 minus cosine theta parenthesis times parenthesis 1 plus cosine theta parenthesis divided by parenthesis 1 minus sine theta parenthesis times parenthesis 1 plus sine theta parenthesis"

OpenStudy (kimberly_pr):

Ok, I will, you're right about that, lol.

OpenStudy (agent0smith):

This just looks awesomely hilarious though lol "parenthesis 1 minus cosine theta parenthesis times parenthesis 1 plus cosine theta parenthesis divided by parenthesis 1 minus sine theta parenthesis times parenthesis 1 plus sine theta parenthesis"

OpenStudy (welshfella):

OK can you expand the top and bottom of this fraction?

OpenStudy (kimberly_pr):

Lol, yeah, idk why I didn't put the pic in the first place XD Oh btw I fanned you @agent0smith and @welshfella I had already fanned you a while back ago. Idk if you guys care about that stuff or not, lol

OpenStudy (kimberly_pr):

Ummm, let me see, give me a sec

OpenStudy (welshfella):

yeah - its nice to know you are appreciated.

OpenStudy (kimberly_pr):

Wouldn't the -cos theta and the +cos theta cancel out? And same with the sin at the bottom? And yes, you both are vey appreciated :)

OpenStudy (kimberly_pr):

I feel like the answer would be cos theta/ sin theta... Am I right?

OpenStudy (welshfella):

you are on the right track this is an example of the 'difference of two squares' (a + b)( a - b) = a^2 + ab - ab - b^2 = a^2 - b^2

OpenStudy (welshfella):

now if we compare that with (1 + cos x)(1 - cos x) we see that it expands to 1 - cos^2 x follow?

OpenStudy (kimberly_pr):

Yes, I follow

OpenStudy (welshfella):

OK in a similar fashion the bottom becomes 1 - sin^2 x so we have 1 - cos^2 x ---------- 1 - sin^2 x which can be simplifies some more.

OpenStudy (welshfella):

do you know the identity sin^2 x + cos^2 x = 1 ?

OpenStudy (kimberly_pr):

Hmmmm, I'm trying to remember... I think the answer would still be cos theta/ sin theta Either that or tan^2 theta, because I know it can't be sin^2 theta or cos^2 theta

OpenStudy (welshfella):

sin^2 x + cos^2 x = 1 so 1 - cos^2 x = sin^2 x and 1 - sin^2 x = cos^2 x make the substitutions

OpenStudy (kimberly_pr):

So it would be sin^2 x/ cos^2x ... Now I'm confused. Sorryyy, this is NOT my forte

OpenStudy (welshfella):

yes that is right now use the identity sin x / cos x = tan x

OpenStudy (kimberly_pr):

So it's tan^2 theta! Right? :)

OpenStudy (welshfella):

to do these problems you need a good knowledge of trig identities.

OpenStudy (welshfella):

YES

OpenStudy (kimberly_pr):

Yayyyy, thanks!!

OpenStudy (welshfella):

yw gtg My daughter needs picking up at the station.

OpenStudy (welshfella):

I'm late already

OpenStudy (kimberly_pr):

Alright, well bye, thanks for everything. Go pick up your daughter and have fun :) I've always wanted a daughter :) I have a baby sister who is 2 years old. She is 14 years younger than me, lol